In a pendulum, kinetic energy is 0 at the extreme positions, so how is the speed maximum at the extremes?????? Pls explain in refernce to this statement!!!!

Grade:9

4 Answers

vikas askiitian expert
13 years ago
dear shivam , let bob of pendulam is raised to heigth H , then it is allowed to fall .... this point will be extrimity of its motion ...now bob will move from this position to lowermost position (mean position) ... from enegy conservation (KE)i + (PE)i = (KE)f + (PE)f (i , f are initial & final state) initially bob was at rest so (KE)i = 0 (at extreme pos) finally bob was moving so (KE)f = K (at mean pos) initially it was at height H so PE stored = mgH (at extreme pos) finally it comes down so its PE becomes 0 .. (at mean pos) since energy is conserved so we can say that all of its PE is converted into its KE when it is at lowermost point(mean pos)... therefore maximum velocity of bob is at mean position ... methametically , v = AWcoswt ..........1 & x = Asinwt ........2 V is max when coswt = + or -1 wt = 0 or pi at 0 or pi , x = 0 x = 0 , means mean position ... therefore max velocity at mean pos only ...
iit jee
13 years ago

hi

potential energy is maximum at the extremes .when it again starts its motion towards the mean position then the potential

energy is converted into kinetic energy.thus the velocity is maximum at the mean position not at the extremes.at the extreme

its acceleration is maximum.

Ritvik Gautam
13 years ago

Actually, at the extreme positions, the speed is not maximum, but the acceleration is. So, due to that acceleration, the speed gets maximum.

 

At the exterme point,

v=0

KE=0

PE=max

a=max

 

At the mean position

v=max

KE=max

PE=0

a=0

 

 

 

Please Approve!!

ajinkya bhole
13 years ago

......Undecided

speed iz max at bottom most part...